[Brahms-l] Shift report 20050220 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 20 2005 - 23:46:38 EST
Shifter leader was EBJ
Shifters were 


1.  The shift begain with beam for about an hour, and then they worked on the beam for 3-4 hours before restoring the beam for data production.

2. I collected some data with the FS at 8deg B2442 and the MRS at 40 A1050.  
I also took some calibration runs.  

3. I only had one problem with the system.  The DAQ crashed (see the elog), but we recovered quickly and began taking data again.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 20 23:47:07 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 20 2005 - 23:47:20 EST